You buy some calla lilies and peonies at a flower store. Calal lillies cost $3.50 each and peonies cost $5.50 each the total cost of 12 flowers is $52. How many calla lilies and how many peonies did you buy?
I tried setting the problem up and solving:
12 (x-5.50)=52

Answer by stanbon(75887)

You buy some calla lilies and peonies at a flower store. Calal lillies cost $3.50 each and peonies cost $5.50 each the total cost of 12 flowers is $52. How many calla lilies and how many peonies did you buy?
----
Quantity: P + C = 12
Value:: 5.5P+3.5C = 52
----
Modify:
35C + 35P = 35*12
35C + 55P = 520
------
20C = 100
P = 5 (# of Peonies)
-----
Solve for "C":
C + P = 12
C + 5 = 12
C = 7 (# of Cala lilies)
=========================
Cheers,
Stan H.